OpenStudy (anonymous):

Which equations show that the set of whole numbers is not closed under subtraction? Choose all answers that are correct. A. 1 – (–2) = 3 B. 1 – 2 = –1 C. 2 – 0 = 2 D. 2 – 4 = –2

zepdrix (zepdrix):

You want to `start with` two elements in W, perform your subtraction, and end up with something outside of that set. In option A, \[\Large\rm 1-(\color{red}{-2})=3\]Notice we're performing our subtraction with an alement from outside of W, so that doesn't show us the property failing subtraction, it's not setup correctly.

OpenStudy (aum):

Whole numbers are: 0, 1, 2, 3, 4, ..... Which equations show that the set of whole numbers is not closed under subtraction? A) 1 - (-2) = 3 1 is a whole number but -2 is not. B) 1 - 2 = -1 1 and 2 are whole numbers. After subtraction, the result is -1 which is NOT a whole number. Therefore, this equation shows that the set of whole numbers is not closed under subtraction. Can you try C and D?

OpenStudy (anonymous):

thanks guys